The economy is characterized by the following equations.
1. Goods Market Equilibrium:
Goods market equilibrium:
Consumption: Investment:
Y = Z
Z=C+I
C = 500 + 0,8Yd, I = 500 + 0,1Y
with Yd = Y

a) Solve for the equilibrium output, consumption and investment. Show your results in an appropriate graph.
b) Assume output is equal to 9000. Compute the overall savings in the economy. Are they equal to investment? Explain!

I know the answers just not the way to solve it. Thank you very much in advance
1
Expert's answer
2019-07-31T10:58:41-0400

a) in equilibrium Y = C + I, so:

Y = 500 + 0.8Y + 500 + 0.1Y,

Y = 10,000.

C = 500 + 0.8×10,000 = 8500,

I = 500 + 0.1×10,000 = 1,500.

b) If output is equal to 9000, then the overall savings equal to investment and they are:

S = I = 500 + 0.1×9,000 = 1,400.